# Practical multi-fidelity machine learning: fusion of deterministic and Bayesian models

### v1

@@ Abstract
Multi-fidelity machine learning methods address the accuracy-efficiency trade-off by integrating scarce, resource-intensive high-fidelity data with abundant but less accurate low-fidelity data. We propose a practical multi-fidelity strategy for problems spanning low- and high-dimensional domains, integrating a non-probabilistic regression model for the low-fidelity with a Bayesian model for the high-fidelity. The models are trained in a staggered scheme, where the low-fidelity model is transfer-learned to the high-fidelity data and a Bayesian model is trained for the residual. This three-model strategy – deterministic low-fidelity, transfer learning, and Bayesian residual – leads to a prediction that includes uncertainty quantification both for noisy and noiseless multi-fidelity data. The strategy is general and unifies the topic, highlighting the expressivity trade-off between the transfer-learning and Bayesian models (a complex transfer-learning model leads to a simpler Bayesian model, and vice versa). We propose modeling choices for two scenarios, and argue in favor of using a linear transfer-learning model that fuses 1) kernel ridge regression for low-fidelity with Gaussian processes for high-fidelity; or 2) deep neural network for low-fidelity with a Bayesian neural network for high-fidelity. We demonstrate the effectiveness and efficiency of the proposed strategies and contrast them with the state-of-the-art based on various numerical examples. The simplicity of these formulations makes them practical for a broad scope of future engineering applications.
Keywords Bayesian machine learning  $\cdot$ Multi-fidelity modeling  $\cdot$ Uncertainty quantification  $\cdot$ Gaussian process regression
@@ 1 Introduction
Expressive non-probabilistic machine learning methods, such as deep neural networks (DNNs), are able to approximate any nonlinear continuous function, provided enough data [1]. However, they can be prone to over-fitting, especially in the presence of noisy data [2], and are not able to predict uncertainty. Bayesian Machine Learning (BML) methods do not have these limitations [3], although their training requires additional computational effort. In this article, we present a practical multi-fidelity (MF) strategy that combines non-probabilistic low-fidelity (LF) surrogates and Bayesian residual surrogates. We demonstrate the simplicity of the proposed MF strategy and consider two common regression scenarios encountered in engineering practice: (1) low-dimensional problems amenable to models with few hyperparameters and easy training; and (2) high-dimensional problems that require more expressive models and involve additional training effort.
@@
MF data arises naturally in most engineering applications. High-fidelity (HF) data is often associated with costly and time-consuming experiments, or with accurate yet computationally expensive simulations. In contrast, LF data is usually obtained by efficient strategies (experimental [4], analytical [5], or computational [6, 7]) at the expense of accuracy. Therefore, LF data is typically acquired faster, often by several orders of magnitude. Using both datasets can be advantageous to train better models as long as the loss in fidelity is compensated by the presence of more low-fidelity data.
@@
The challenges associated with developing MF regression models arise from the need to define one regression model (a surrogate) per fidelity, and in the subsequent definition of the interaction between these models [8, 9]. Without loss of generality, we focus on MF regression models with only two levels of fidelity. We start by introducing MF regression in a general form and by quickly reviewing common choices for surrogate models. We continue by reviewing common MF strategies, which involve the same type of surrogate model across fidelities. Then, we propose two MF strategies using non-probabilistic LF regression models together with Bayesian residual models.
@@
Summary of contributions. We introduce a general framework for MF regression and demonstrate its versatility. We hypothesize that most MF regression cases in practice can be covered by considering simple linear regression (LR) transfer-learning models and the following two options for low- and high-fidelity models: 1) kernel ridge regression (KRR) and Gaussian process regression (GPR), leading to the KRR-LR-GPR model; or 2) a deep neural network (DNN) and a Bayesian neural network (BNN), leading to the DNN-LR-BNN model. The KRR-LR-GPR model has few parameters and hyperparameters, facilitating training but limiting scalability (limits on data dimension and size, as well as training and inference time). The DNN-LR-BNN model is more flexible and more challenging to train, but it is applicable for low- and high-dimensional data with few scalability limits. The simpler model (KRR-LR-GPR) should be used when possible, while the DNN-LR-BNN model is expected to be applicable to most practical cases. In the limit of having complex correlations between low- and high-fidelity, a DNN-BNN model can be used where the BNN takes care of both transfer-learning and residual modeling, although this leads to the most difficult model to train and is expected to have narrower applicability.
@@ 2 Methodology and related work
For simplicity of notation, we start by considering datasets with $d$-dimensional inputs $\mathbf{x}$ but one-dimensional output $y$. Then, the LF dataset contains $\mathbf{x}^{l}_{n}$ input and $y^{l}_{n}$ output points, where $n=1,...,N^{l}$ are the $N^{l}$ points in this dataset (the superscript $l$ refers to low-fidelity). The HF dataset is equivalently defined by $\mathbf{x}^{h}_{n}$ and $y^{h}_{n}$, where $n=1,...,N^{h}$ are the $N^{h}$ HF points in this dataset. In this article, we define a MF regression model $f^{h}(\mathbf{x})$ as a conjunction of three models: (1) the LF model $f^{l}(\mathbf{x})$ that is trained on LF data $\{\mathbf{x}^{l},y^{l}\}$; (2) a transfer-learning model $g(\mathbf{x})$ that transforms the LF model to the HF data; (3) a residual model $r(\mathbf{x})$ (if necessary) that captures the difference between transfer-learned LF model and the HF data. Therefore, a general description can be formulated as:
@@
In the data-scarce literature [10, 11], MF models use GPRs for both $f^{l}(\mathbf{x})$ and $r(\mathbf{x})$ and consider a linear transfer-learning model $g(\mathbf{x}):=g(f^{l}(\mathbf{x}))=f^{l}(\mathbf{x})\rho$, such that:
@@
Other investigations focus on using a transfer-learning model of higher complexity, for example, Bayesian models [12, 13], while setting $r(\mathbf{x}):=0$, leading to:
@@
Here, we argue in favor of generalizing these modeling assumptions by establishing a simple transfer-learning model $g(\mathbf{x})$ that is trained on HF data, and by choosing an appropriate deterministic low-fidelity model $f^{l}(\mathbf{x})$ and a Bayesian high-fidelity residual $r(\mathbf{x})$.
@@ 2.1 Related work for data-scarce scenarios
In the data-scarce regime, GPR or Kriging [14] stand out due to their elegant derivation and exact integration for Gaussian likelihoods and priors (see Appendix A.2 for a short introduction). Early in their development, GPR was extended to handle MF problems according to Eq. (2), in a method originally called Co-Kriging [15, 16]. This method is based on expanding the covariance function of GPRs in the form of
@@
Different variants of this formulation can be found, including multi-task GPR [11, 17] that consider more than one output, or latent mapping GPR for cases with categorical variables [18, 19]. Unfortunately, Co-Kriging has complexity of order $\mathcal{O}\left((N^{l}+N^{h})^{3}\right)$ due to the inversion involved when calculating the determinants of the covariance matrix $\mathbf{C}$, imposing practical limits on the number of data points and the dimensionality [20, 10, 21, 22]. To this end, Hierarchical Kriging [23] and corresponding variants [24, 25, 17] considering a diagonal sub-matrix of $\mathbf{C}$ have been proposed. For example, Scale Kriging [26, 27] assumes a fixed hyperparameter for the transfer model, i.e. $\rho=1$, while recursive Co-Kriging [28] uses fast cross-validation for identifying parameters. In practice, however, it can be challenging to strike a balance between applying MF-GPR to large datasets (improving complexity) without degrading performance significantly [29, 30]. Furthermore, the literature is scarce on their application to noisy MF data [31].
We also note that there are other methods of interest in data-scarce MF problems. For example, a MF polynomial chaos expansion model [31] that also relies on Eq. (2) has recently been proposed. The method was demonstrated for noisy HF data, but considering noiseless LF data. The scalability of these approaches is mainly limited by the amount of LF data, usually several times more than that of HF data. In contrast, deterministic MF models have shown better scalability at the sacrifice of not predicting uncertainty. Examples include MF linear regression [32], MF Radial Basis Kernel (RBF) regression [33], and MF support vector regression [34]. Yet, the combination of deterministic and Bayesian models for MF regression merits further investigation.
@@ 2.2 Related work for data-rich scenarios
Machine learning models trained on data-rich MF datasets can be seen in different contexts, for instance, assisting design [35, 36], optimization [37], and in uncertainty quantification [38]. Focusing on regression problems, deterministic models such as DNNs have been shown to obtain a reasonable point estimate instead of the full posterior. For example, the MF-DNN approach proposed by Aydin et al. [39] sequentially passes data from the LF to the HF into the same DNN with an error metric for fidelity switching. Transfer learning based on MF dataset [40, 41] uses the LF dataset to pre-train the weights of a DNN and then shifts to the HF dataset to finalize training. Other examples include a MF Graph Neural Network (GNN) developed by Black et al. [42], and a MF-DNN architecture introduced by Motamed et al. [38] that establishes the correlation between the HF and LF models via a neural network. Other strategies have involved the concatenation of LF- and HF-DNNs [43, 44, 12], as well as Recurrent Neural Networks [45], and Convolutional Neural Networks [46].
Although deterministic MF-DNNs are common in the literature, they are ineffective in predicting uncertainty, are prone to over-fitting, and do not provide a principled way to handle noisy data [13, 44, 43]. Instead, BNNs [47] provide a theoretical framework that naturally addresses these limitations but introduces a few other issues. In particular, the integration of the posterior distribution and the posterior predictive distribution (PPD) is intractable analytically. Usually, Bayesian inference approaches such as Markov Chain Monte Carlo (MCMC) [47, 48], Variational Inference (VI) [49], Monte Carlo Dropout [50], among many others [51] are used for integrating the posterior distribution. However, the prohibitive training and inference time limits the adoption of MF-BNNs as opposed to MF-DNNs. Recent work has shown promise, for example, Meng et al. [13] extended their previous concatenated MF-DNN architecture [43] by replacing the HF DNN with a BNN and using Hamiltonian Monte Carlo to perform inference in a reasonable amount of time [52, 53]. Still, its effectiveness has only been demonstrated in low-dimensional examples. In addition, Baptiste et al. [54] proposed to use both Bayesian models for HF and LF datasets, where a GPR was used in the LF and a BNN was used in HF. However, they cautioned that the strategy was only suitable for solving low-dimensional problems because of the time-consuming inference procedure. In summary, considering BNNs in the context of MF data is challenging due to the trade-off between inference time and accuracy. Although in theory, using both BNNs for LF and HF datasets is possible, it has also been shown that the inference time can be worse than a simple combination of a GPR and a BNN [54]. Furthermore, despite considering different architectures, most rely on the linear model of Eq. (2), and they are applied in the data-scarce scenario without considering noisy data, especially at the LF level.
3 Practical multi-fidelity Bayesian machine learning (MF-BML)
@@
In practice, creating a MF model involves so many choices [10, 11] that it is difficult to perform principled model selection and hyperparameter optimization, especially for large datasets. We propose a simple strategy summarized in Eq. (2) to develop practical MF models based on two ideas, as shown in Fig. 1. First, consider a deterministic LF model $f^{l}(\mathbf{x})$ together with a probabilistic model of the residual $r(\mathbf{x})$. Second, consider a transfer-learning model $g(\mathbf{x})$ as a simple linear regression whose basis function is the LF model. As we demonstrate in the remainder of the paper, the proposed strategy leverages the advantages of deterministic and probabilistic models while keeping the transfer learning model simple.
Figure 1: Illustration of how the practical MF-BML evolves from GPR. (a) GPR with explicit basis function, in which a first linear regression is executed based on the basis function $\mathbf{h(x)}=[1,\mathbf{x},\mathbf{x}^{2},...]^{T}$ and where $\delta(\mathbf{x})$ is assumed to be a zero mean GPR for characterizing the residual. (b) the practical MF-BML follows the same concept by utilizing the LF surrogate as the basis. A linear transfer-learning function on $f^{l}(\mathbf{x})$ adjusts the prediction for the HF, and facilitates the determination of the residual $r(\mathbf{x})$ by a Bayesian model with a simple prior.
Thus, the transfer-learning model $g(\mathbf{x})$ in Eq. (2) is proposed to be linear (in the weights) but considering its basis functions as the LF model. In the case of considering only one LF model, we write the proposed MF-BML model as follows:
$f^{h}(\mathbf{x})$ $=g\left(f^{l}(\mathbf{x})\right)+r(\mathbf{x})$ (4)
$=\bm{m}(\mathbf{x})^{T}\bm{\rho}+r(\mathbf{x})$ (5)
@@
where $\mathbf{m(x)}=[1,f^{l}(\mathbf{x}),f^{l}(\mathbf{x})^{2},\ldots,f^{l}(\mathbf{x})^{M-1}]^{T}$ is the vector with $M$ polynomial basis formed with the LF model, and $\bm{\rho}=[\rho_{0},\rho_{1},\ldots,\rho_{M-1}]^{T}$ is the corresponding coefficient vector to be determined from training on the HF data, i.e. these coefficients are not imposed as hyperparameters. As previously mentioned in Section 2.1, a majority of MF models simplify the transfer-learning model to become linear, i.e. they reduce Eq. (4) to Eq. (2) by considering $\mathbf{m(x)}=f^{l}(\mathbf{x})$ and then considering $\rho_{1}$ as a hyperparameter.
In addition to defining the transfer-learning model, there are many possible choices for LF and HF models. Table 1 summarizes common models, as reviewed in the introduction. We suggest two different MF model choices for spanning a large number of data-scarce and data-rich MF regression problems: 1) kernel ridge regression together with Gaussian process regression (KRR-LR-GPR) [^1] Naming convention for the MF model (¡LF¿-¡transfer-learning¿¡residual¿): the abbreviation on the left is the surrogate for the LF model, the middle is the transfer-learning model, the last abbreviation is the model for the residual. In the absence of one of the models, for example not considering the residual, then the abbreviation only has two models (e.g. DNN-BNN refers to a deep neural network model for LF with a BNN doing the transfer-learning and no residual). for data-scarce or low-dimensionality scenarios; and 2) deep neural network together with a Bayesian neural network (DNN-LR-BNN) for data-rich scenarios.
@@
[c] LF surrogates HF surrogates Linear regression (LR) Gaussian process regression (GPR) Kernel ridge regression (KRR) Polynomial chaos expansion (PCE) Deep neural network (DNN) Bayesian neural network (BNN) Support Vector Regression (SVR) Gaussian process regression (GPR)* Polynomial chaos expansion (PCE)*
@@
There is an advantage in choosing a deterministic LF model because it only acts as a trend function that is transformed by a transfer-learning function, leading to $f^{h}(\mathbf{x})-r(\mathbf{x})$. Such LF models can be efficiently trained due to their lower complexity (better scalability) when compared to probabilistic models, handling large LF datasets with or without noise, but without characterizing uncertainty.
@@
Choosing a probabilistic model that complements the deterministic LF model, it is possible to make HF prediction that include uncertainty quantification while harnessing the typical advantages of probabilistic models: less prone to overfitting and better approximation quality even in the presence of few HF data.

### v2

@@ Abstract
Multi-fidelity machine learning methods address the accuracy-efficiency trade-off by integrating scarce, resource-intensive high-fidelity data with abundant but less accurate low-fidelity data. We propose a practical multi-fidelity strategy for problems spanning low- and high-dimensional domains, integrating a non-probabilistic regression model for the low-fidelity with a Bayesian model for the high-fidelity. The models are trained in a staggered scheme, where the low-fidelity model is transfer-learned to the high-fidelity data and a Bayesian model is trained to learn the residual between the data and the transfer-learned model. This three-model strategy – deterministic low-fidelity, transfer-learning, and Bayesian residual – leads to a prediction that includes uncertainty quantification for noisy and noiseless multi-fidelity data. The strategy is general and unifies the topic, highlighting the expressivity trade-off between the transfer-learning and Bayesian models (a complex transfer-learning model leads to a simpler Bayesian model, and vice versa). We propose modeling choices for two scenarios, and argue in favor of using a linear transfer-learning model that fuses 1) kernel ridge regression for low-fidelity with Gaussian processes for high-fidelity; or 2) deep neural network for low-fidelity with a Bayesian neural network for high-fidelity. We demonstrate the effectiveness and efficiency of the proposed strategies and contrast them with the state-of-the-art based on various numerical examples and two engineering problems. The results indicate that the proposed approach achieves comparable performance in both mean and uncertainty estimation while significantly reducing training time for machine learning modeling in data-scarce scenarios. Moreover, in data-rich settings, it outperforms other multi-fidelity architectures by effectively mitigating overfitting.
Keywords Bayesian machine learning  $\cdot$ Multi-fidelity modeling  $\cdot$ Uncertainty quantification  $\cdot$ Gaussian process regression  $\cdot$ Bayesian Neural Network
List of Abbreviations
  • DNN – Deep Neural Network
  • BML – Bayesian Machine Learning
  • MF – Multi-Fidelity
  • LF – Low-Fidelity
  • HF – High-Fidelity
  • LR - Linear Regression
  • KRR – Kernel Ridge Regression
  • GPR – Gaussian Process Regression
  • KRR-LR-GPR - The proposed practical multi-fidelity Bayesian machine learning model for the data-scarce scenario, which uses Kernel Ridge Regression as the low-fidelity surrogate, Linear Regression model for transfer-learning, and Gaussian process regression to learn for the residual.
  • BNN – Bayesian Neural Network
  • MF-BML - Multi-Fidelity Bayesian Machine Learning
  • DNN-LR-BNN - The proposed practical MF Bayesian machine learning model for the data-rich scenario, which uses Deep Neural Network as the low-fidelity surrogate, Linear Regression model for transfer-learning, and Bayesian Neural Network to learn for the residual.
  • DNN-BNN - Multi-fidelity Bayesian Neural Network employing low-fidelity deep neural network and Bayesian neural network as transfer-learning model [1].
  • RBF – Radial Basis Kernel
  • PCE – Polynomial Chaos Expansion
  • SGLD – Stochastic Gradient Langevin Dynamics
  • pSGLD – preconditioned Stochastic Gradient Langevin Dynamics
  • NRMSE – Normalized Root Mean Square Error
  • TLL – Test Log-Likelihood
  • PPD – Posterior Predictive Distribution
  • QoI - Quantity of Interest
  • SVE - Stochastic Volume Element
  • FEA - Finite Element Analysis
  • SCA - Self-consistent Clustering Analysis
@@ 1 Introduction
Expressive deterministic (non-probabilistic) machine learning methods, such as DNNs, can approximate any nonlinear continuous function, providing enough data [2]. However, they can be prone to over-fitting, especially in the presence of noisy data [3], and are not able to predict uncertainty. BML methods do not have these limitations [4], although their training requires additional computational effort. In this article, we present a practical MF strategy that combines non-probabilistic LF surrogates and Bayesian residual surrogates. We demonstrate the simplicity of the proposed MF strategy and consider two common regression scenarios encountered in engineering practice: (1) low-dimensional problems amenable to models with few hyperparameters and easy training; and (2) high-dimensional problems that require more expressive models and involve additional training effort.
@@
MF data arises naturally in most engineering applications. HF data is often associated with costly and time-consuming experiments, or with accurate yet computationally expensive simulations. In contrast, LF data is usually obtained by efficient strategies (experimental [5], analytical [6], or computational [7, 8]) at the expense of accuracy. Therefore, LF data is typically acquired faster, often by several orders of magnitude. Using both datasets can be advantageous to train better models as long as the loss in fidelity is compensated by the presence of more LF data.
@@
The challenges associated with developing MF regression models arise from the need to define one regression model (a surrogate) per fidelity, and in the subsequent definition of the interaction between these models [9, 10]. Without loss of generality, we focus on MF regression models with only two levels of fidelity. We start by introducing MF regression in a general form and by reviewing common choices for surrogate models. We continue by reviewing common MF strategies employing the same type of surrogate model across fidelity levels. Finally, we propose two MF strategies using non-probabilistic LF regression models together with Bayesian residual models.
@@
Summary of contributions. We introduce a practical framework for MF regression and demonstrate its versatility. We hypothesize that most MF regression cases in practice can be covered by considering simple LR transfer-learning models and with one of the following two options for LF and HF models: 1) using KRR and GPR, leading to the KRR-LR-GPR model; or 2) using DNN and BNN, leading to the DNN-LR-BNN model. The KRR-LR-GPR model has few parameters and hyperparameters, facilitating training but limiting scalability (limits on data dimension and size, as well as training and inference time). The DNN-LR-BNN model is more flexible and more challenging to train but is applicable to low- and high-dimensional data with few scalability limits. The simpler model (KRR-LR-GPR) should be used when possible. In limited cases exhibiting complex correlations between LF and HF data, both transfer-learning and residual learning can be achieved simultaneously by a BNN – denoted as DNN-BNN models [1]. However, DNN-BNN models are more difficult to train and are expected to have narrower applicability.
@@ 2 Methodology and related work
For simplicity of notation, we start by considering datasets with $d$-dimensional inputs $\mathbf{x}$ but one-dimensional output $y$. Then, the LF dataset contains $\mathbf{x}^{l}_{n}$ input and $y^{l}_{n}$ output points, where $n=1,...,N^{l}$ are the $N^{l}$ points in this dataset (the superscript $l$ refers to LF). The HF dataset is equivalently defined by $\mathbf{x}^{h}_{n}$ and $y^{h}_{n}$, where $n=1,...,N^{h}$ are the $N^{h}$ HF points in this dataset. In this article, we propose to define an MF regression model $f^{h}(\mathbf{x})$ as a combination of three models: (1) the LF model $f^{l}(\mathbf{x})$ that is trained on LF data $\{\mathbf{x}^{l},y^{l}\}$; (2) a transfer-learning model $g(\mathbf{x})$ that transforms the LF model to the HF data; and (3) a residual model $r(\mathbf{x})$ (if necessary) that captures the difference between transfer-learned LF model and the HF data. Therefore, a general description can be formulated as:
@@
In the data-scarce literature [11, 12], MF models use GPRs for both $f^{l}(\mathbf{x})$ and $r(\mathbf{x})$ and implicitly assume a linear transfer-learning model $g(\mathbf{x}):=g(f^{l}(\mathbf{x}))=f^{l}(\mathbf{x})\rho$, such that:
@@
Other investigations use more expressive transfer-learning models without defining a different model for the residual [1, 13]:
@@
Here, we argue in favor of generalizing these modeling assumptions by establishing a simple transfer-learning model $g(\mathbf{x})$ that is trained on HF data, choosing an appropriate deterministic LF model $f^{l}(\mathbf{x})$ and a Bayesian HF model for the residual $r(\mathbf{x})$.
@@ 2.1 Related work for data-scarce scenarios
In the data-scarce regime, GPR or Kriging [14] stand out due to their elegant derivation and exact integration for Gaussian likelihoods and priors –see Section A.2 for a short introduction. Early in their development, GPR was extended to handle MF problems according to Equation 2, in a method originally called Co-Kriging [15, 16]. This method is based on expanding the covariance function of GPRs in the form of
@@
Different variants of this formulation can be found, including multi-task GPR [12, 17] that consider more than one output, or latent mapping GPR for cases with categorical variables [18, 19]. Unfortunately, Co-Kriging has a complexity of order $\mathcal{O}\left((N^{l}+N^{h})^{3}\right)$ due to the inversion of the covariance matrix $\mathbf{C}$ (see Equation 18), imposing practical limits on the dimensionality and number of training data points that can be considered [20, 11, 21]. In part this issue is mitigated in Hierarchical Kriging [22] and its variants [23, 24, 17] by defining a diagonal sub-matrix of $\mathbf{C}$. Recursive Co-Kriging [25] employs a fast hyperparameter optimization strategy and cross-validation to identify the parameter $\rho$, leading to comparable performance to Co-Kriging in some cases. In addition, Scale Kriging [26, 27] adopts a simplification by considering a fixed parameter with $\rho=1$ in Equation 2 for the transfer-learning model. However, in practice, striking a balance between training accurate MF-GPR models and improving their complexity such that they are trained on large datasets remains a significant challenge [28, 29]. Furthermore, the literature is scarce on their application to noisy MF data [30].
In addition, data-scarce MF problems can be addressed by other methods. For example, MF-PCE [30] has been applied to noisy HF data and noiseless LF data. This method also relies on Equation 2, but suffers from similar scalability issues as MF-GPR. In contrast, deterministic MF models have shown better scalability but do not predict uncertainty. Examples include MF linear regression [31], MF-RBF regression [32], and MF support vector regression [33]. Inspired by the strengths of deterministic and Bayesian methods, this article explores the combination of deterministic and Bayesian models in MF regression.
@@ 2.2 Related work for data-rich scenarios
Machine learning models trained on data-rich MF datasets can be seen in different contexts, including design [34, 35], optimization [36], and uncertainty quantification [37]. When solving regression problems with deterministic models such as DNNs, predicting the full posterior is avoided by finding a point estimate. For example, the MF-DNN approach proposed by Aydin et al. [38] passes data sequentially from the LF to the HF into the same DNN and includes an error metric for fidelity switching. Other strategies involve transfer-learning [39, 40] by pre-training the weights of a DNN on the LF dataset and then finalizing training on the HF dataset. Other relevant examples include an MF Graph Neural Network developed by Black et al. [41], and an MF-DNN architecture introduced by Motamed et al. [37] that establishes the correlation between the HF and LF models via another neural network. Some strategies concatenate LF- and HF-DNN models [42, 43, 13], as well as Recurrent Neural Networks [44], and Convolutional Neural Networks [45].
Although deterministic MF-DNNs are common in the literature, they are ineffective in predicting uncertainty and are prone to over-fitting [1, 43, 42]. BNNs [46] address these limitations but introduce other issues. In particular, the integration of the posterior distribution and the posteriod predictive distribution (PPD) is intractable analytically, so Bayesian inference is performed by Markov Chain Monte Carlo [46, 47], Variational Inference [48], Monte Carlo Dropout [49], among others [50]. Unfortunately, Bayesian inference is computationally expensive and makes it challenging to adopt these models for large or high-dimensional datasets. This is visible in the promising work of Meng et al. [1], who extended a deterministic MF-DNN architecture [42] by replacing the HF-DNN with a BNN, concatenating it with a LF-DNN, and using Hamiltonian Monte Carlo to perform inference [51, 52]. Yet, the method is only demonstrated on small datasets. Similar limitations are reported using Bayesian models for both LF and HF data, as in the work of Baptiste et al. [53] where GPR is used for LF and a BNN for HF. The same authors note that using BNNs for both LF and HF datasets is possible, but the inference time becomes worse than combining GPR with BNN.
The literature spans different architectural choices involving deterministic or Bayesian models. Most strategies rely on the linear model shown in Equation 2, and apply methods to data-scarce scenarios without considering noisy data, especially at the LF level. To our knowledge, addressing the trade-off between inference time and accuracy remains a knowledge gap in the literature.
3 Practical multi-fidelity Bayesian machine learning
@@
In practice, creating MF models involves so many choices [11, 12] that is difficult to perform principled model selection and hyperparameter optimization, especially for large datasets. We propose a simple strategy summarized in Equation 1 to develop practical MF models based on two ideas, as shown in Figure 1. First, consider a deterministic LF model $f^{l}(\mathbf{x})$ together with a probabilistic model of the residual $r(\mathbf{x})$. Second, consider a transfer-learning model $g(\mathbf{x})$ as a simple linear regression model whose features become the outputs of the LF model, i.e. $g\left(f^{l}(\mathbf{x})\right)$. As we demonstrate in the remainder of the paper, the proposed strategy leverages the advantages of deterministic and probabilistic models while keeping the transfer-learning model simple.
Figure 1: Schematic (a) shows a linear regression model with basis function $\mathbf{h(x)}=[1,\mathbf{x},\mathbf{x}^{2},...]^{T}$ and coefficients $\boldsymbol{\beta}$ that is augmented by the residual $\delta(\mathbf{x})$ modeled by zero mean GPR. Schematic (b) shows the proposed MF-BML strategy where the transfer-learning model $g\left(f^{l}(\mathbf{x})\right)$ is a linear regression model whose features are the outputs of the LF surrogate model $f^{l}(\mathbf{x})$ obtained by training on LF data. The linear transfer-learning function that acts on $f^{l}(\mathbf{x})$ adjusts the LF model to the HF data by determining the coefficients $\boldsymbol{\rho}$, and facilitates the determination of the residual $r(\mathbf{x})$ by a Bayesian model with a simple prior.
Writing the transfer-learning model $g(\mathbf{x})$ in Equation 2 as a linear model (linear in the weights) and considering its features as the LF model leads to:
$f^{h}(\mathbf{x})$ $=g\left(f^{l}(\mathbf{x})\right)+r(\mathbf{x})$
$=\bm{m}(\mathbf{x})^{T}\bm{\rho}+r(\mathbf{x})$
@@
where $\mathbf{m(x)}=[1,f^{l}(\mathbf{x}),f^{l}(\mathbf{x})^{2},\ldots,f^{l}(\mathbf{x})^{M-1}]^{T}$ is the vector with $M$ polynomial basis features of the LF model, and $\bm{\rho}=[\rho_{0},\rho_{1},\ldots,\rho_{M-1}]^{T}$ is the corresponding coefficient vector to be determined from training on the HF data, i.e. these coefficients are not imposed as hyperparameters. This contrasts with the models reviewed in Section 2.1 that reduce Equation 4 to Equation 2 by considering $\mathbf{m(x)}=f^{l}(\mathbf{x})$ and then considering $\rho_{1}$ as a hyperparameter.
In addition to defining the transfer-learning model, there are many possible choices for LF and HF models. Table 1 summarizes common models, as reviewed in the introduction. We suggest two different MF model choices for spanning a large number of data-scarce and data-rich MF regression problems: 1) kernel ridge regression together with Gaussian process regression (KRR-LR-GPR) [^1] Naming convention for the MF model (¡LF model¿-¡transfer-learning model¿-¡residual model¿): the abbreviation on the left is the surrogate for the LF model, the middle is the transfer-learning model, and the last abbreviation is the model for the residual. In the absence of one of the models, for example not considering the residual, then the abbreviation only has two models (e.g. DNN-BNN refers to a DNN model for LF with a BNN doing the transfer-learning and no residual). for data-scarce or low-dimensionality scenarios; and 2) DNN together with a Bayesian neural network (DNN-LR-BNN) for data-rich or multi-output scenarios.
@@
[c] LF surrogates HF surrogates Linear regression (LR) Gaussian process regression (GPR) Kernel ridge regression (KRR) Polynomial chaos expansion (PCE) Deep neural network (DNN) Bayesian neural network (BNN) Support Vector Regression (SVM) Gaussian process regression (GPR) * Polynomial chaos expansion (PCE) *
@@
LF datasets are usually larger than HF datasets because, by definition, LF data is acquired faster than HF data. If the LF dataset is large, deterministic LF models are advantageous because probabilistic models are not sufficiently scalable. In addition, deterministic LF models have lower complexity and fewer hyperparameters, while still being able to handle noisy data (aleatoric uncertainty).
@@
The key disadvantage of choosing deterministic LF models is their inability to characterize epistemic (or model) uncertainty (at the LF). However, this will be shown to have limited relevance for the LF model when the residual or the transfer-learning model is probabilistic. In such cases, the HF predictions include uncertainty quantification and harness the typical advantages of probabilistic models.